# JEL classification: C33, C38, C55

### v1

@@ 2.3 Sparse Factor Processes
In view of the discussion in Section 2.2, the traditional APCA estimators are difficult to interpret, especially when the dimension is large. In the past decades, various sparse PCA or sparse factor modeling techniques have been proposed in the statistical and econometrics literature. See, for example, Zou, Hastie, and Tibshirani (2006), Witten, Tibshirani, and Hastie (2009), and Kristensen (2017), among other. Sparse PCA restricts the cardinality of the weight vectors for the PCs so that the PCs are sparse linear combinations of the underlying cross-sectional variables. By setting the factor weights to zero, sparse PCA facilitates the interpretation of the PCs, which improves interpretability without sacrificing the explanatory ability of the PCs or factors. To the best of our knowledge, all the aforementioned approaches assume that the loadings are sparse in the population model, and formulate the estimation of sparse loadings as a regularized optimization problem to estimate principle components with a penalty term, which allows for developing asymptotic inferential theory. Nevertheless, the assumption of sparse population loadings may not be satisfied in many datasets. For example, in finance, the exposure of asset returns to a market factor is universal and non-sparse. In addition, according to Camacho, Smilde, Saccenti, and Westerhuis (2020), the sparse PCA method presents disparate and limited performance when modeling sparse data, even in the optimistic case of noise-free data.
@@
To illustrate the interpretations of the proposed approach, we take the panel of asset returns in finance as an example. First of all, the factor process measures the overall co-movement of all stock returns over the time horizon considered in the data set. When there is no co-movement over some certain period, the factors would be zero at such time points. Under such circumstances, the stock returns would be driven by their idiosyncratic terms because there are not enough government policies, influential news, or other major events that can drive all the cross-sectional movement, and consequently, the systematic response might be not significant. Second, each loading vector is a linear combination of the corresponding series over the time horizon, which measures the average dependence on the factors for the given time period of the data. When the factors are zero at a specific time point, the contribution of the series at such time point to the loading would be negligible.
2.4 Examples of Sparse Factors in Time Horizons
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
    Monitoring and Forecasting. Stock and Watson (1998) and Forni, Hallin, Lippi, Reichlin, et al. (2000) have demonstrated that factor models provide an effective way to monitor economic activities, as large-dimensional economic panel data can be described by a small number of factors. In other words, business cycles can be modeled by the co-movements of economic variables, which correspond to the common factors ${\mathbf{f}}_{t}$’s in Model (2). When the factors are sparse, we can monitor economic activities according to the systematic risks detected by our approach. Additionally, Stock and Watson (1998) also considered forecasting inflation with the common factors (diffusion indices) extracted from a large panel of macroeconomic variables and found that these factors can improve forecasting accuracy.
@@ 2.5 Sparse APCA: Formulation and Estimation
We first introduce some further notation used in the estimation procedure. For an $N\times 1$ vector ${\mathbf{v}}=(v_{1},...,v_{N})^{\prime}$, $\|{\mathbf{v}}\|_{2}=\sqrt{\sum_{i=1}^{2}v_{i}^{2}}$ is the Euclidean norm and $\|{\mathbf{v}}\|_{0}=\text{card}\{\text{support}({\mathbf{v}})\}$ is the cardinality or number of non-zero elements in ${\mathbf{v}}$. Let $\mathbb{V}$ be the set of $N\times r$ orthogonal matrices and $\mathbb{V}^{\perp}$ be the set consisting of the orthogonal complements of the ones in $\mathbb{V}$. For a matrix ${\mathbf{V}}=({\mathbf{v}}_{1},...,{\mathbf{v}}_{r})\in\mathbb{V}$ and a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$, $\|{\mathbf{V}}\|_{0}\leq{\mathbf{s}}$ implies that $\|{\mathbf{v}}_{i}\|_{0}\leq s_{i}$ for $1\leq i\leq r$.
@@ 2.5.1 One-Factor Case
which is a non-convex problem in general. In fact, it is NP-hard because it can be reduced to the subset selection problem for ordinary least-squares problem; see Moghaddam, Weiss, and Avidan (2006).
In fact, the optimization in (6) is still a sparse PCA problem symbolically, and numerous methods have been developed to obtain sparse eigenvectors during the past decades. See the regularization method with elastic net in Zou, Hastie, and Tibshirani (2006) and the penalized matrix decomposition (PMD) algorithm using $\ell_{1}$-penalty in Witten, Tibshirani, and Hastie (2009), among others.
@@
where ${\mathbf{S}}$ is the matrix obtained by the noisy observation ${\mathbf{x}}_{t}$, $\boldsymbol{\Sigma}$ is a symmetric matrix whose eigenvectors are the true ones, and ${\mathbf{E}}$ is a random perturbation. The decomposition of (8) is formulated in (IA.1) of the online Appendix. If the true largest eigenvector ${\mathbf{v}}_{1}$ of $\boldsymbol{\Sigma}$ is sparse, then it is natural to recover ${\mathbf{v}}_{1}$ from the noisy matrix ${\mathbf{S}}$. This recovery is guaranteed when the error ${\mathbf{E}}$ is of a smaller order using the well-known $\sin\theta$ theorem in Davis and Kahan (1970) under the approximate factor model. Second, for any given vector ${\mathbf{u}}_{0}\in R^{N}$, the power method estimates the first eigenvector by
@@ 2.5.2 Multi-Factor Case
In this section, we consider the general case when there are multiple factor processes in Model (3), i.e. $r>1$. In the presence of more than one common factor process, we need to apply Algorithm 1 multiple times and extract all the common factors in a sequential way. Specifically, when $\widehat{\mathbf{v}}_{1}$ is given via Algorithm 1, we subtract the projection on first factor component of the panel and the residual would be $\widetilde{\mathbf{X}}={\mathbf{X}}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{X}}=({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{X}}$. The resulting scaled covariance $\widetilde{\mathbf{S}}_{1}=({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime})$. Then the second eigenvector $\widehat{\mathbf{v}}_{2}$ can be obtained by solving the following optimization problem (11) below:
$\widetilde{\mathbf{v}}_{2}=\arg\max{\mathbf{v}}^{\prime}\widetilde{\mathbf{S}}_{1}{\mathbf{v}},\,\,\text{subject to}\,\,{\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq s_{2},$ (11)
and the second eigenvector is $\widehat{\mathbf{v}}_{2}=\widetilde{\mathbf{v}}_{2}/\|\widetilde{\mathbf{v}}_{2}\|_{2}$. Consequently, the second estimated factor process is $\underline{\widehat{\mathbf{f}}}_{2}=\sqrt{T}\widehat{\mathbf{v}}_{2}$. We use the normalization condition ${\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1$ to maximize the additional variance of the original matrix ${\mathbf{S}}$, which is the same as the deflation framework in Mackey (2008). In other words, we need to eliminate the effect of the previous eigenvector directions by a projection method. Consequently, we formulate the procedure in Algorithm 2 below.
@@
From the optimization problem in (11) and the procedure in Algorithm 2, we cannot enforce the orthogonality and sparsity at the same time, which is similar to the case in the sparse PCA framework. When there is no $\ell_{0}$-constraint in (6) and (11), we can easily obtain that $\widehat{\mathbf{v}}_{1}$ and $\widehat{\mathbf{v}}_{2}$ are the two normalized and orthogonal eigenvectors associated with the top two eigenvalues of ${\mathbf{S}}$ and, the approach reduces to the traditional APCA method. One of the differences between the one-factor case in (6) and the multi-factor case in (11) is that the estimated factors and loadings in the one-factor case satisfy the conditions in (4), whereas those obtained in the multi-factor case may not. It is important to note that this discrepancy is only a finite sample outcome. Asymptotically, we can demonstrate that the multi-factor case still satisfies the identification conditions in (4). Similar outcomes are also observed in the sparse PCA literature, including Zou, Hastie, and Tibshirani (2006) and Witten, Tibshirani, and Hastie (2009), as well as in the sparse factor modeling frameworks discussed in Kristensen (2017).
@@ 2.7 Choice of the Tuning Parameters
In practice, the cardinality of each column in ${\mathbf{F}}$ is unknown, and we may choose them by the cross-validation method which is commonly used in machine learning literature. In fact, even though the sparsity parameters $s_{1},...,s_{r}$ of $r$ factor processes may not be the same, it is still a convenient way to assume that $s_{1}=...=s_{r}=s_{0}$ which can simplify the cross-validation procedure significantly. According to the theoretical results in Section 3 and the proofs in the Appendix, when the $r$ sparsity parameters are distinct, the convergence of the estimator for $s^{*}$ can be achieved, where $s^{*}=\max(s_{1},...,s_{r})$. Therefore, we assume that the sparsity parameters of each column are the same in this section. The cross-validation procedure is as follows.
@@
$PC(s)=R(s,\widetilde{\mathbf{F}}_{1}^{s})+rsg(N_{1},T).$
@@
$\widehat{s}=\arg\min_{s\in\mathbb{S}}PC(s)=\arg\min_{s\in\mathbb{S}}\{R(s,\widetilde{\mathbf{F}}_{1}^{s})+rsg(N_{1},T)\},$ (15)
implying that we choose the optimal $s$ which results in minimal testing errors in the cross-validations. There are many choices for the penalty function $g(N_{1},T)$ as shown in Theorem 4 in Section 3 below. We will discuss the required properties of $g(N_{1},T)$ such that we can establish the consistency of $\widehat{s}$ in theory.
In practice, we may replace the candidate set $\mathbb{S}$ by $1\leq s\leq smax$ where $smax$ is a prescribed bounded integer. Note that the cross-validation method is different from the classical one in the machine learning literature where the samples are partitioned along the time horizon, whereas we partition the samples over the cross-section in this paper. In fact, we can even perform more cross-validation experiments in order to obtain an optimal parameter $s$. For example, we can choose a sufficiently large integer $J>0$, and we can obtain $J$ testing errors as that in (14) for each fixed $s$ by partitioning the samples for $J$ times. We may denote the testing error by $R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s})$ for the $j$-th random partition and the average testing error for $s$ is defined as
@@ 3 Theoretical Properties
In this section, we present the asymptotic consistency and stochastic bounds of the estimators from Section 2 under under the framework where $N,T\rightarrow\infty$. Assuming the number of factors $r$ is fixed, and $s_{i}$ can be either fixed or slowly increasing, we first derive the asymptotic results under known $r$ and $s_{i}$. We also demonstrate the consistency of the estimators for $r$ and $s_{i}$ thereafter. The mathematical proofs are provided in the Appendix.
@@
Assume $E\|{\mathbf{f}}_{t}\|_{2}^{4}\leq C$ and the covariance satisfies $\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}={\mathbf{I}}_{r}$.
@@
There exists a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$ such that $\|{\mathbf{F}}\|_{0}\leq{\mathbf{s}}$, where ${\mathbf{F}}$ is defined as that in (3).
@@
Assumption 1 is standard to characterize the dynamic dependence of the factor processes. See, for example, the Lemma 1 in the appendix of Gao, Ma, Wang, and Yao (2019). Assumptions 2-3 impose an identification condition and a sparsity structure on the true factor process ${\mathbf{F}}$. Assumption 4 specifies that the top eigenvalues of the covariance of the panel data are distinct in order to avoid multiplicity of eigenvalues. Assumptions 2 and 4 provide identifiability for the Model (3) so that the factors and the loading matrix can be uniquely determined asymptotically. See Bai and Ng (2013) for a detailed argument. Assumption 5 is only made to simplify the theoretical derivations. It can be relaxed to a weaker assumption such as Conditions (c.i)–(c.iii) in Bai and Ng (2013). Similar assumptions are also made in Huang, Jiang, Li, Tong, and Zhou (2022) to simplify the theoretical derivation. Assumption 7 is used to derive the Bernstein-type concentration inequality in establishing the convergence rates of the estimated factors, and it can be relaxed to much weaker ones as those in Merlevède, Peligrad, and Rio (2011).
@@
Let Assumptions 17 in the Appendix hold and $\widehat{\mathbf{v}}_{1}$ be the first solution of (6). As $N,T\rightarrow\infty$, it holds that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right),$
@@
From Theorem 1, we see that the angle between the estimated direction $\widehat{\mathbf{v}}_{1}$ and the true one ${\mathbf{v}}_{1}$ is asymptotically equal to zero or $\pi$ if $s_{1}\log(N)/(NT)\rightarrow 0$. In fact, we can rewrite $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}$ as $\cos(\theta)$ where $\theta$ is the angle between $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$. Some remarks for Theorem 1 are as follows
@@
$\sqrt{1-\cos^{2}(\theta)}=|\sin(\theta)|=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right).$
If $s_{1}\log(N)/(NT)\rightarrow 0$, we have that $\sin(\theta)\rightarrow 0$ asymptotically, implying that $\theta\rightarrow 0$ or $\theta\rightarrow\pi$. Therefore, $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$ coincide on the same line.
@@
$\|\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}\|_{2}=\sqrt{2(1-\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})}\leq\sqrt{2(1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2})}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{NT}}+\frac{s_{1}\log(N)}{NT}+\frac{1}{T}\right),$
@@
$\|\underline{\widehat{\mathbf{f}}}_{1}-\underline{{\mathbf{f}}}_{1}\|_{2}=O_{p}\left(\sqrt{\frac{s_{1}\log(N)}{N}}+\frac{s_{1}\log(N)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
implying that we can uniquely determine the factors asymptotically if $s_{1}\log(N)=o(N)$ as $N,T\rightarrow\infty$.
Next, we present the theorem concerning the consistency of all the estimated factors along the time horizon. For the distance between two matrices ${\mathbf{H}}_{1}$ and ${\mathbf{H}}_{2}$, there are several measures that can be used in the literature. For example, We adopt the discrepancy measure used by Pan and Yao (2008): for two $T\times r$ semi-orthogonal matrices ${\bf H}_{1}$ and ${\bf H}_{2}$ satisfying the condition ${\bf H}_{1}^{\prime}{\bf H}_{1}={\bf H}_{2}^{\prime}{\bf H}_{2}={\mathbf{I}}_{r}$, the difference between the two linear spaces $\mathcal{M}({\bf H}_{1})$ and $\mathcal{M}({\bf H}_{2})$ is measured by

### v2

@@ 2.3 Sparse Factor Processes
In view of the discussion in Section 2.2, the traditional APCA estimators are difficult to interpret, especially when the dimension is large. In the past decades, various sparse PCA or sparse factor modeling techniques have been proposed in the statistical and econometrics literature. See, for example, Zou et al. (2006), Witten et al. (2009), and Kristensen (2017), among other. Sparse PCA restricts the cardinality of the weight vectors for the PCs so that the PCs are sparse linear combinations of the underlying cross-sectional variables. By setting the factor weights to zero, sparse PCA facilitates the interpretation of the PCs, which improves interpretability without sacrificing the explanatory ability of the PCs or factors. To the best of our knowledge, all the aforementioned approaches assume that the loadings are sparse in the population model, and formulate the estimation of sparse loadings as a regularized optimization problem to estimate principle components with a penalty term, which allows for developing asymptotic inferential theory. Nevertheless, the assumption of sparse population loadings may not be satisfied in many datasets. For example, in finance, the exposure of asset returns to a market factor is universal and non-sparse. In addition, according to Camacho et al. (2020), the sparse PCA method presents disparate and limited performance when modeling sparse data, even in the optimistic case of noise-free data.
@@ 2.4 Sparse APCA: Formulation and Estimation
We first introduce some further notation used in the estimation procedure. For an $N\times 1$ vector ${\mathbf{v}}=(v_{1},...,v_{N})^{\prime}$, $\|{\mathbf{v}}\|_{2}=\sqrt{\sum_{i=1}^{2}v_{i}^{2}}$ is the Euclidean norm and $\|{\mathbf{v}}\|_{0}=\text{card}\{\text{support}({\mathbf{v}})\}$ is the cardinality or number of non-zero elements in ${\mathbf{v}}$. Let $\mathbb{V}$ be the set of $N\times r$ semi-orthogonal matrices and $\mathbb{V}^{\perp}$ be the set consisting of the orthogonal complements of the ones in $\mathbb{V}$. For a matrix ${\mathbf{V}}=({\mathbf{v}}_{1},...,{\mathbf{v}}_{r})\in\mathbb{V}$ and a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$, $\|{\mathbf{V}}\|_{0}\leq{\mathbf{s}}$ implies that $\|{\mathbf{v}}_{i}\|_{0}\leq s_{i}$ for $1\leq i\leq r$.
@@ 2.4.1 One-Factor Case
which is a non-convex problem in general. In fact, it is NP-hard because it can be reduced to the subset selection problem for ordinary least-squares problem; see Moghaddam et al. (2006).
In fact, the optimization in (6) is still a sparse PCA problem symbolically, and numerous methods have been developed to obtain sparse eigenvectors during the past decades. See the regularization method with elastic net in Zou et al. (2006) and the penalized matrix decomposition (PMD) algorithm using $\ell_{1}$-penalty in Witten et al. (2009), among others.
@@
where ${\mathbf{S}}$ is the matrix obtained by the noisy observation ${\mathbf{x}}_{t}$, $\boldsymbol{\Sigma}$ is a symmetric matrix whose eigenvectors are the true ones, and ${\mathbf{E}}$ is a random perturbation. The decomposition of (8) is formulated in (S.1) of the online Appendix. If the true largest eigenvector ${\mathbf{v}}_{1}$ of $\boldsymbol{\Sigma}$ is sparse, then it is natural to recover ${\mathbf{v}}_{1}$ from the noisy matrix ${\mathbf{S}}$. This recovery is guaranteed when the error ${\mathbf{E}}$ is of a smaller order using the well-known $\sin\theta$ theorem in Davis and Kahan (1970) under the approximate factor model. Second, for any given vector ${\mathbf{u}}_{0}\in R^{N}$, the power method estimates the first eigenvector by
@@
Algorithm 1 can consistently estimate sparse eigenvectors under a factor structure when the corresponding eigenvalues diverge. However, this property may not extend to general covariance matrices. For instance, it is possible to construct a diagonal covariance matrix with sparse eigenvectors, where the algorithm may fail to correctly identify the sparse eigenvectors associated with top eigenvalues for some initial vectors. For instance, consider ${\mathbf{S}}=\mbox{diag}(2,1)$. In this case, the leading eigenvector is ${\mathbf{v}}_{1}=(1,0)^{\prime}$, but the algorithm will incorrectly output ${\mathbf{v}}_{1}=(0,1)^{\prime}$ if the initial vector is ${\mathbf{u}}_{0}=(1,3)^{\prime}$. This occurs because the top eigenvalue is not sufficiently large. However, if ${\mathbf{S}}=\mbox{diag}(4,1)$, where the top eigenvalue is significantly larger—as is typically the case in a factor model—the algorithm correctly identifies the leading eigenvector.
@@ 2.4.2 Multi-Factor Case
In this section, we consider the general case when there are multiple factor processes in Model (3), i.e. $r>1$. In the presence of more than one common factor process, we need to apply Algorithm 1 multiple times and extract all the common factors in a sequential way. Specifically, when $\widehat{\mathbf{v}}_{1}$ is given via Algorithm 1, we subtract the projection on first factor component of the panel and the residual would be $\widetilde{\mathbf{X}}={\mathbf{X}}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{X}}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{X}}$. The resulting scaled covariance $\widetilde{\mathbf{S}}_{1}=({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime})$. Then the second eigenvector $\widehat{\mathbf{v}}_{2}$ can be obtained by solving the following optimization problem (11) below:
$\widetilde{\mathbf{v}}_{2}=\arg\max{\mathbf{v}}^{\prime}\widetilde{\mathbf{S}}_{1}{\mathbf{v}},\,\,\text{subject to}\,\,{\mathbf{v}}^{\prime}({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq s_{2},$ (11)
and the second eigenvector is $\widehat{\mathbf{v}}_{2}=\widetilde{\mathbf{v}}_{2}/\|\widetilde{\mathbf{v}}_{2}\|_{2}$. Consequently, the second estimated factor process is $\underline{\widehat{\mathbf{f}}}_{2}=\sqrt{T}\widehat{\mathbf{v}}_{2}$. We use the normalization condition ${\mathbf{v}}^{\prime}({\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}=1$ to maximize the additional variance of the original matrix ${\mathbf{S}}$, which is the same as the deflation framework in Mackey (2008). In other words, we need to eliminate the effect of the previous eigenvector directions by a projection method. Consequently, we formulate the procedure in Algorithm 2 below.
@@
From the optimization problem in (11) and the procedure in Algorithm 2, we cannot enforce the orthogonality and sparsity at the same time, which is similar to the case in the sparse PCA framework. When there is no $\ell_{0}$-constraint in (6) and (11), we can easily obtain that $\widehat{\mathbf{v}}_{1}$ and $\widehat{\mathbf{v}}_{2}$ are the two normalized and orthogonal eigenvectors associated with the top two eigenvalues of ${\mathbf{S}}$ and, the approach reduces to the traditional APCA method. One of the differences between the one-factor case in (6) and the multi-factor case in (11) is that the estimated factors and loadings in the one-factor case satisfy the conditions in (4), whereas those obtained in the multi-factor case may not. It is important to note that this discrepancy is only a finite sample outcome. Asymptotically, we can demonstrate that the multi-factor case still satisfies the identification conditions in (4). Similar outcomes are also observed in the sparse PCA literature, including Zou et al. (2006) and Witten et al. (2009), as well as in the sparse factor modeling frameworks discussed in Kristensen (2017).
@@ 2.6 Choice of the Tuning Parameters
In practice, the cardinality of each column in ${\mathbf{F}}$ is unknown, and we may choose them by the cross-validation method which is commonly used in machine learning literature. In fact, even though the sparsity parameters $s_{1},...,s_{r}$ of $r$ factor processes may not be the same, it is still a convenient way to assume that $s_{1}=...=s_{r}=s_{0}\asymp T^{\alpha}$ which can simplify the cross-validation procedure significantly. According to the theoretical results in Section 3 and the proofs in the Appendix, when the $r$ sparsity parameters are distinct, the convergence of the estimator for $s^{*}$ can be achieved, where $s^{*}=\max(s_{1},...,s_{r})$. Therefore, we assume that the sparsity parameters of each column are the same in this section. The cross-validation procedure is as follows.
@@
$PC(s)=R(s,\widetilde{\mathbf{F}}_{1}^{s})+r\frac{s}{T^{\alpha}}g(N_{1},T),$
@@
$\widehat{s}=\arg\min_{s\in\mathbb{S}}PC(s)=\arg\min_{s\in\mathbb{S}}\{R(s,\widetilde{\mathbf{F}}_{1}^{s})+r\frac{s}{T^{\alpha}}g(N_{1},T)\},$ (15)
implying that we choose the optimal $s$ which results in minimal testing errors in the cross-validations. Note that the sparsity strength parameter $\alpha$is unknown and must be specified prior to applying the criterion. Its value depends on the specific application and involves an element of subjective judgment. Importantly, $\alpha$ does not influence the study of risk factors over time, as it only reflects the relative strengths of the nonzero factors. In practice, we recommend setting $\alpha=1/2$, which is sufficient for identifying a robust set of systematic risk events and factors in empirical analyses.
There are many choices for the penalty function $g(N_{1},T)$ as shown in Theorem 4 in Section 3 below. We will discuss the required properties of $g(N_{1},T)$ such that we can establish the consistency of $\widehat{s}$ in theory.
In practice, we may replace the candidate set $\mathbb{S}$ by $smin\leq s\leq smax$ where $smin$ and $smax$ are some prescribed integers. Note that the cross-validation method is different from the classical one in the machine learning literature where the samples are partitioned along the time horizon, whereas we partition the samples over the cross-section in this paper. In fact, we can even perform more cross-validation experiments in order to obtain an optimal parameter $s$. For example, we can choose a sufficiently large integer $J>0$, and we can obtain $J$ testing errors as that in (14) for each fixed $s$ by partitioning the samples for $J$ times. We may denote the testing error by $R_{j}(s,\widetilde{\mathbf{F}}_{1,j}^{s})$ for the $j$-th random partition and the average testing error for $s$ is defined as
@@ 3 Theoretical Properties
In this section, we present the consistency and asymptotic bounds of the estimators from Section 2 under the framework where $N,T\rightarrow\infty$. Assuming the number of factors $r$ is fixed, and $s_{i}$ can be either fixed or slowly increasing, we first derive the asymptotic results under known $r$ and $s_{i}$. We also demonstrate the consistency of the estimators for $r$ and $s_{i}$ thereafter. The mathematical proofs are provided in the Appendix.
@@
Assume $\frac{1}{T}\sum_{t=1}^{T}{\mathbf{f}}_{t}{\mathbf{f}}_{t}^{\prime}\rightarrow_{p}{\mathbf{I}}_{r}$.
@@
There exists a vector ${\mathbf{s}}=(s_{1},...,s_{r})^{\prime}$ with $s_{i}\asymp T^{\alpha_{i}}$ such that $\|{\mathbf{F}}\|_{0}\leq{\mathbf{s}}$, where ${\mathbf{F}}$ is defined as that in (3) and $\alpha_{i}\in(0,1]$ for $1\leq i\leq r$.
@@
Assumption 1 is standard to characterize the dynamic dependence of the factor processes. See, for example, the Lemma 1 in the appendix of Gao et al. (2019). Assumptions 2-3 establish an identification condition and a sparsity structure for the true factor process ${\mathbf{F}}$. While the factors are sparse, Assumption 2 is not restrictive, as it is reasonable to assume $s_{i}$ diverges and Assumption 2 naturally holds under Assumption 3 by setting $\alpha_{i}=1$ and specifying $s_{i}/T=c_{i}\in(0,1)$, which is in line with the setting in random matrix theory. Note that we could adopt the framework in Uematsu and Yamagata (2022), which treats factors as weak with varying strengths. However, specifying these strengths in a given application can be challenging. Instead, we follow the strong factor framework in Kristensen (2017) under sparsity, and the strength of nonzero factors may diverge if $\alpha_{i}<1$ in Assumption 3. This divergence is merely a matter of the data’s magnitude and does not affect our goal of identifying nonzero factors over time. Assumption 4 specifies that the top eigenvalues of the covariance of the panel data are distinct in order to avoid multiplicity of eigenvalues. Assumptions 2 and 4 provide identifiability for the Model (3) so that the factors and the loading matrix can be uniquely determined asymptotically. See Bai and Ng (2013) for a detailed argument.Assumption 5 is primarily introduced to simplify the theoretical derivations. It can be relaxed to weaker conditions, such as Conditions (c.i)–(c.iii) in Bai and Ng (2013), or replaced with strongly mixing conditions in both spatial and temporal dimensions. Similar independence assumptions have also been adopted in Huang et al. (2022) to streamline theoretical analyses. Assumption 7 is used to derive the Bernstein-type concentration inequality in establishing the convergence rates of the estimated factors, and it can be relaxed to much weaker ones as those in Merlevède et al. (2011).
@@
Let Assumptions 17 hold and $\widehat{\mathbf{v}}_{1}$ be the first solution of (6). As $N,T\rightarrow\infty$, it holds that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{NT}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right),$
@@
To maintain clear interpretation, we continue to use $s_{1}$ instead of $s_{1}\asymp T^{\alpha_{1}}$ in the theorem. From Theorem 1, we see that the angle between the estimated direction $\widehat{\mathbf{v}}_{1}$ and the true one ${\mathbf{v}}_{1}$ is asymptotically equal to zero or $\pi$ if $s_{1}\log(T)/(NT)\asymp\log(T)/(NT^{1-\alpha_{1}})\rightarrow 0$. In fact, we can rewrite $\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1}$ as $\cos(\theta)$ where $\theta$ is the angle between $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$. Some remarks for Theorem 1 are as follows.
@@
$\sqrt{1-\cos^{2}(\theta)}=|\sin(\theta)|=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{NT}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right).$
If $s_{1}\log(T)/(NT)\rightarrow 0$, we have that $\sin(\theta)\rightarrow 0$ asymptotically, implying that $\theta\rightarrow 0$ or $\theta\rightarrow\pi$. Therefore, $\widehat{\mathbf{v}}_{1}$ and ${\mathbf{v}}_{1}$ coincide on the same line.
@@
$\|\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}\|_{2}=\sqrt{2(1-\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})}\leq\sqrt{2(1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2})}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{NT}}+\frac{s_{1}\log(T)}{NT}+\frac{1}{T}\right),$
@@
$\|\underline{\widehat{\mathbf{f}}}_{1}-\underline{{\mathbf{f}}}_{1}\|_{2}=O_{p}\left(\sqrt{\frac{s_{1}\log(T)}{N}}+\frac{s_{1}\log(T)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
implying that we can uniquely determine the factors asymptotically if $s_{1}\log(T)=o(N)$ as $N,T\rightarrow\infty$.
(iii) When the sparsity is proportional to the sample size, i.e., $s_{1}\asymp T$, we have that
$\frac{1}{\sqrt{T}}\|\underline{\widehat{\mathbf{f}}}_{1}-\underline{{\mathbf{f}}}_{1}\|_{2}=O_{p}\left(\sqrt{\frac{\log(T)}{N}}+\frac{1}{T}\right),$
which is in line with the conventional result in factor modeling if we ignore the $\log(T)$ term. See Gao and Tsay (2024) for details.
Next, we present the theorem concerning the consistency of all the estimated factors along the time horizon. For the distance between two matrices ${\mathbf{H}}_{1}$ and ${\mathbf{H}}_{2}$, there are several measures that are used in the literature. For example, we adopt the discrepancy measure used by Pan and Yao (2008): for two $T\times r$ semi-orthogonal matrices ${\bf H}_{1}$ and ${\bf H}_{2}$ satisfying the condition ${\bf H}_{1}^{\prime}{\bf H}_{1}={\bf H}_{2}^{\prime}{\bf H}_{2}={\mathbf{I}}_{r}$, the difference between the two linear spaces $\mathcal{M}({\bf H}_{1})$ and $\mathcal{M}({\bf H}_{2})$ is measured by